A car of mass m travelling at speed v crashes into the rear of a truck of mass 2m that is at rest and in neutral at an intersection. If the collision is perfectly inelastic, what is the speed of the combined car and truck after the collision?


1
Expert's answer
2021-02-15T00:58:48-0500
"mv=(m+2m)v_c,""v_c=\dfrac{mv}{(m+2m)}=\dfrac{mv}{3m}=\dfrac{1}{3}v."